A data set of size n = 25 is presented by the frequency table below. Only part of the information is Given. Write the full raw data, i.e., list all the values in this data set. (The order does not matter).

0pnR0b1JnqUAAAAAElFTkSuQmCC
Answer #1

Distinct
Values

Cumulative
Counts

Percents
1 3 3/25 = 12%
3 7 (7-3)/25 = 16%
6 7 + 25*0.4 = 17 40%
10 19 (19-17)/25 = 8%
15 25 (25-19)/25 = 24%

Raw data = 1, 1, 1, 3, 3, 3, 3, 6, 6, 6, 6, 6, 6, 6, 6, 6, 6, 10, 10, 15, 15, 15, 15, 15, 15